# On the Operational Resilience of CBDC: Threats and Prospects of Formal Validation for Offline Payments

### v1

@@ Abstract
Information and communication technologies are by now employed in most activities, including economics and finance. Despite the extraordinary power of modern computers and the vast amount of memory, some results of theoretical computer science imply the impossibility of certifying software quality in general.
With the exception of safety-critical systems, this has primarily concerned the information processed by confined systems, with limited socio-economic consequences. In the emerging era of technologies for exchanging digital money and tokenized assets over the Internet – such as central bank digital currencies (CBDCs) – even a minor bug could trigger a financial collapse.
Although the aforementioned impossibility results cannot be overcome in an absolute sense, there exist formal methods that can provide assertions of computing systems correctness. We advocate their use to validate the operational resilience of software infrastructures enabling CBDCs, with special emphasis on offline payments as they constitute a very critical issue.
@@ 1.1 CBDC and Digital Euro: A Digital Form of Cash
Central bank digital currencies (CBDCs) are set to become the latest technological step in the long history of money. After barter, cowrie shells, metal coins, and banknotes, the 20th century saw the introduction of digital money through the immateriality of bits stored on computers operated by banks, with account balances managed via bank transfers and debit or credit card transactions. In our modern developed societies, where most of our activities are digitally mapped, the way people pay has a digital dimension too, with more and more consumers around the world increasingly preferring to pay by using digital electronic devices rather than traditional physical cash, such as banknotes or coins. Therefore, in a context where the use of physical cash is declining, the idea of a CBDC directly controlled by the central bank has gained growing attention; it would represent the next step forward for legal tender issued by a central bank.
@@
A CBDC would be a digital form of cash that complements banknotes and coins, giving people an additional choice of payment modalities. It would be freely available to all citizens and issued by the central bank. If managed under a federalist model, commercial banks would play a key role as the front-end interface with customers. Under this paradigm, commercial banks would remain responsible for managing the relationship with end customers, activating the necessary know-your-customer/anti-money-laundering (KYC/AML) procedures, distributing approved CBDC wallets to customers, and resolving any issue that might arise. However, citizens would be free to receive their salaries, pensions, and other types of small incoming payments directly into their own wallets. At the same time, they would be able to spend their money by using an online or offline service provided by the central bank, which would be secure and guarantee the appropriate level of privacy.
@@
It is well known that today 137 countries and currency unions, representing 98% of global GDP, are exploring CBDC solutions through ongoing investigations, studies, models, and pilot projects on the topic [5].
@@
In particular, the European Central Bank (ECB) plans to introduce a digital euro [30] as a retail and wholesale CBDC. It will be freely available and free of charge for any digital payment to all citizens in the euro area. After an Investigation Phase that took place from October 2021 to October 2023, aimed at assessing feasibility, benefits, design options, possible risks, and issues, a two-year Preparation Phase was launched in November 2023. The objective is to test the potential issuance, define the necessary rules, identify the infrastructure required for smooth and secure usage, and establish the legal frameworks and operational rulebook under which to pursue the issuance. Another key element is conducting user research to ensure that the digital euro project under development is able to address the vast majority of the needs of European citizens and is inclusive from the very beginning in all its forms.
@@
From the legal point of view, in 2023 the European Commission has approved a proposal for its regulation [30].
@@ 1.2 The Challenge of Operational Resilience in CBDC
Computing systems, along with software methodologies and programming languages needed to design and implement them, have faced a growing complexity. From sequential systems, in which a single operation at a time can be executed, the first evolution has been toward concurrent and distributed systems [24], in which multiple computing devices run simultaneously and respectively exchange information through a shared memory or via message passing. The second evolution, started with the financial crisis of 2008 and the consequent decreasing trust in institutional intermediaries, has led to decentralized systems [60], where parties not knowing each other perform transactions by relying on a distributed ledger and a consensus mechanism that ensure trust in an algorithmic way.
Several advanced paradigms are by now in place. They range from mobile computing, where devices are no longer tied to physical locations (IoT – Internet of Things [57]), global computing, which implements the abstraction of a single computer accessible anywhere anytime (cloud, edge, fog infrastructures [25, 17]), and autonomic computing, typical of systems capable of adapting to unpredictable changes (via sensors, actuators, policies [51]), to machine learning, deep learning, and large language models (AI – Artificial Intelligence [68, 80, 48]). Moreover, ongoing research is focusing on future paradigms such as quantum computing [61], which should make computation faster, and reversible computing [52], which should reduce energy consumption.
@@
This high complexity exacerbate the problem of software quality; the interested reader is referred to [81] for a list of major software disasters and their consequences in terms of costs, human lives, and environmental impact. Software quality is not only about code functional correctness, because it is equally important to avoid poor performance, security breaches, and bad interfaces. Moreover, the usually adopted approach of testing software is not enough as it does not guarantee the absence of errors. Therefore, it is of paramount importance to adopt software verification techniques, which are mostly based on the early development of software models and hence need to be accompanied by model-driven software engineering methodologies.
@@
The trend of cybercrimes involving hackers stealing money is closely monitored by specialized institutions. For example, the International Monetary Fund (IMF) reports more than 20,000 cyber incidents in the traditional financial sector between 2004 and 2023, with total losses of 10 billion USD (see Figure 3.2 of [47]). In cryptocurrencies and the decentralized finance (DeFi) sector, supported by public blockchains, Chainalysis reports nearly 14 billion USD in losses from 1,329 hacking events between 2015 and 2024 [18]. Even if not all the breaches described above are necessarily linked to structural flaws in the software, the phenomenon remains deeply concerning.
@@ 1.3 Formal Methods for Ensuring Software Quality of CBDC
To overcome these two highly undesirable situations, one might consider building two test programs. The first one, $T_{l}$, is designed so that, when supplied with the code lines of $P$, it can answer “Yes” or “No” in finite time to the question: “Does $P$ enter an infinite loop in some situations?”. The second one, $T_{f}$, is able to answer “Yes” or “No” in finite time to the question: “Does $P$ compute the function $f_{P}$?”. Unfortunately, two fundamental results of computability theory – due to Turing [77] and Rice [67] – establish the impossibility of constructing such tests for all programs. The practical consequence of these two results is that upon every software release there remains a residual risk of malfunction.
@@
On the specification side, formal models of software can be provided at different levels of abstraction and revolve around automata, algebras, and logics. An automaton [45] consists of a set of states and a set of transitions between states. Every state of an automaton represents a state of the computation, like for instance the current statement to be executed in a program along with the value contained in every program variable, while a transition describes a state change, which can be accompanied by possible inputs and outputs. There is a hierarchy of automata classes, ranging from finite-state automata to Turing machines, which correspond to the hierarchy of formal languages established by Chomsky [19]. Moreover, in addition to traditional automata in which states are global, there are variants called Petri nets [65, 66] in which the underlying graph is bipartite so that states can be represented in a distributed way across loci of computations, which is well suited for concurrent and distributed software.
@@
Process algebras [58, 43, 39, 6, 15] are more abstract than automata because they provide a number of operators whereby to obtain complex system models by combining simpler ones. Among these operators we mention sequential, alternative, and parallel compositions, which allow one to express the fact that two processes – intended as the behaviors of two systems – are executed one after the other or alternative to each other or run in parallel, respectively. An important notion in this setting is that of behavioral equivalence, with bisimilarity – intended as the capability of mimicking each other’s behavior stepwise – being one of the most important approaches, which identifies syntactically different processes that feature the same observable behavior. Compositional modeling is accompanied by compositional reasoning when the considered equivalence is a congruence – i.e., it allows to replace equals with equals like in arithmetic – with respect to the process algebraic operators.
@@
Logics are even more abstract in that they express the behavioral properties that a computing system should possess. The simplest example is given by Hoare logic [42, 22] and its variants, including separation logic [63], which are used to enrich programs with annotations containing logical formulas that state properties that should be valid in those points of programs execution. Other examples are modal and temporal logics [46, 33]. They respectively are extensions of classical logic with modal or temporal operators. The former denote the possibility or the necessity of performing certain activities in a given order. The latter express a property to be valid in the next step or the fact that a certain property has to hold until another one becomes satisfied, along at least one computation or all computations.
@@
On the verification side, we mention in particular model checking [56, 21, 7] and equivalence checking [15]. A model checker takes as inputs a program model, described as an automaton or a process term, and a property it should possess, formalized as a modal or temporal logic formula, then communicates whether the program model meets that property or not. In case of failure, it provides diagnostic information in the form of a program model execution that does not fulfill the property. An equivalence checker considers instead a model of the specification of the program and a model of the implementation of the program, then checks whether the two models are behaviorally equivalent to each other or not. In case of failure, it exhibits a distinguishing modal or temporal logic formula, i.e., a formula that is satisfied by only one of the two models.
@@
The aforementioned modeling and verification techniques, originally developed for functional features of programs, have been subsequently extended to deal with software architectures and quantitative aspects too [3, 38, 71, 2, 41, 40, 1, 23, 36]. In this way, it is possible to address also systems including probabilistic behaviors – think, e.g., of cryptocurrencies based on proof-of-stake consensus – or real-time constraints – think, e.g., of the deadline by which a bank transfer can be canceled. Moreover, they have been adapted to address security properties as well, like for instance the use of equivalence checking to assess noninterference of information flows among different security levels [35, 32, 26]. Many software tools have been developed – and employed in large-scale case studies – to implement formal modeling and verification, like for instance CADP [34], mCRL2 [36], GreatSPN [2], PEPA Workbench [41], TwoTowers [3], Spin [44], NuSMV [20], Uppaal [54], PRISM [50], Modest Toolset [37].
@@ 2 Issues and Risks of Offline Payments in a CBDC
Another important advantage of the offline payments solution is that it would enable the financial inclusion of unbanked people, who could use offline payments as a substitute for cash. This could also address the consequences of the declining use of cash [12], with the CBDC acting as a medium for pseudonymous or anonymous payments [27]. Furthermore, in a scenario in which cash is gradually abandoned as a medium for transferring value, commercial banks might find it economically disadvantageous to offer services to unbanked people, who are de facto financially excluded groups. Therefore, the introduction of offline payments as an instrument of financial inclusion is becoming a necessity to support the most vulnerable parts of our societies, which would otherwise be excluded from the economy. To achieve this, only a minimal technological infrastructure would be required – essentially, a smartphone, a basic computer, or a laptop.
In a report by the World Bank Group, it is estimated that globally 1.4 billion adults are unbanked [82]. As for the EU, a 2024 Economic Bulletin of the ECB [29] states that “nearly one out of five adults (19.4%) in the euro area reports not having either debit or credit cards or payment accounts”. As a consequence, the offline payments solution with cash-like features is strongly recommended as an instrument for financial inclusion also in the context of the future digital euro. Indeed, it constitutes the second requirement in the “Report on a digital euro” by the ECB [28]:
@@
A further element supporting the offline payments solution derives from a survey conducted by the Innovation Hub of the Bank for International Settlements (BIS). It shows that “49% of central banks surveyed consider offline payments with retail CBDC to be vital, while another 49% deemed it to be advantageous” [8].
Unfortunately, even if highly recommended, the offline payments solution presents some challenges and issues from operational and technical points of view. In an expository study by the Bank of Finland, a case study named “Project Pluto” is discussed, aimed at studying the risks of offline CBDCs [62]. In contrast to the Riksbank’s e-krona pilot in Sweden [73, 74, 75], the Pluto application focused on the token-based model and utilizes reusable tokens.
In a note of the Bank of Canada [59] it is explicitly written that:
@@
The Bank of England also aimed to assess whether it is technically feasible to implement a secure offline payment functionality for a digital pound. The conclusions highlight some critical issues [12]:
@@
But the most detailed study on the secure feasibility of the offline payments solution is the one conducted by BIS, which has recently produced a series of analytical documents to illustrate “Project Polaris”. The first report, entitled “A Handbook for Offline Payments with CBDC” [8], outlines the current technical solutions for offline payments with a CBDC, the possible risks, threats, and vulnerabilities associated with them, and the countermeasures to mitigate these issues. The second report, “A Security and Resilience Framework for CBDC Systems” [9], is devoted to a study of the objectives and design criteria required to achieve security and resilience in a CBDC framework. The third report, “Closing the CBDC Cyber Threat Modeling Gaps” [10], addresses the problem of analyzing the adequacy of current cybersecurity standards and frameworks through real-world observations of security flaws; the goal is to develop cyber threat models that need to be addressed to achieve full control over the security necessary in a CBDC framework. The fourth and final report, entitled “A High-Level Design Guide for Offline Payments with CBDC” [11], outlines the suggested design choices for offline payment solutions, also analyzing the security, risk, and impact on resilience of the offline system.
@@
Generally speaking, as noted in [14], we must state that basing part of the security of offline solutions on the assumption of inviolability of secure elements [^1] A secure element is a microprocessor chip with enhanced security protections that should prevent unauthorized access. It can be embedded in smart cards and smartphones or in removable subscriber identity module (SIM) cards. is simply naive. Furthermore, in general even a single compromised device could print and spend counterfeit money without restrictions. For example, a hacker able to break a single secure element could send counterfeit money to a colluding device. Moreover, without adequate countermeasures, offline payment recipients cannot distinguish legitimate payments from counterfeit, so central banks would be forced to bear these risks of fraud.
We briefly illustrate here a tentative and non-exhaustive list of threats based in part on the first report of “Project Polaris” [8] by the BIS and on the expository study of the Bank of Finland [62]:
@@
The physical devices used in the framework of offline payments are digital-electronic devices that can leak electromagnetic fields while operating. It is therefore possible, for a bad actor, to analyze the variations of these fields and perform signal analysis on the electromagnetic radiation. This could allow them to deduce some information about the electrical quantities that physically represent the bit strings associated with the digital representation of the CBDC tokens or the private key involved in a transaction.
@@
The double-spending issue can be associated with several previously described situations or contexts. Moreover, note that the double-spending issue for offline CBDC payments is different from the more famous problem of double-spending cryptocurrencies on a public and disintermediated blockchain. In the latter case, the problem of consensus on an asynchronous network must be solved, but it is well known that the extended version of the Byzantine generals problem does not admit theoretical solutions [31]. Nevertheless, practical solutions based on proof-of-work [60] and proof-of-stake [49] protocols offer some level of security. In contrast, in the case of offline payments for a CBDC, trust needs to be ensured by the tamper-resistant nature of the user device and the shared cryptographic protocols of the wallet. It is therefore necessary to ensure that double-spending attempts will be rejected by the payee’s wallet.
@@
Figure 1: The risk of devices staying offline: the bordered ones always remain offline [62].
@@
This is probably the most insidious problem that, to the best of our knowledge, has not yet been solved. Suppose that there are two chains of offline transactions, #1 and #2, as in Figure 1. In both transaction chains, a certain amount X of tokens is moved, while the bordered smartphones always remain offline [62].

### v2

@@ Abstract
Information and communication technologies are by now employed in most activities, including economics and finance. Despite the extraordinary power of modern computers and the vast amount of available memory, some results of theoretical computer science imply the impossibility of certifying software quality in general.
With the exception of safety-critical systems, this has primarily concerned the information processed by confined systems, with limited socio-economic consequences. In the emerging era of technologies for exchanging digital money and tokenized assets over the Internet, such as in particular central bank digital currencies (CBDCs), even a minor bug could trigger a financial collapse.
Although the aforementioned impossibility results cannot be overcome in an absolute sense, there exist formal methods that can provide correctness assertions for computing systems. We advocate their use to validate the operational resilience of software infrastructures enabling CBDCs, with special emphasis on offline payments as they constitute a very critical issue.
@@ 1.1 CBDC and Digital Euro: A Digital Form of Cash
Central bank digital currencies (CBDCs) are set to become the latest technological step in the long history of money. After barter, cowrie shells, metal coins, and banknotes, the 20th century saw the introduction of digital money through the immateriality of bits stored on computers operated by banks, with account balances managed via bank transfers and debit or credit card transactions. In our modern developed societies, where most of our activities are digitally mapped, the way people pay has a digital dimension too, with more and more consumers around the world increasingly preferring to pay by using digital electronic devices rather than traditional physical cash, such as banknotes or coins. Therefore, in a context where the use of physical cash is declining, the idea of a CBDC directly controlled by the central bank has gained growing attention; it would represent the next step forward for legal tender issued by the central bank.
@@
A CBDC would be a digital form of cash that complements banknotes and coins, giving people an additional choice of payment modalities. It would be issued by the central bank and freely available to all citizens. If managed under a federalist model, commercial banks would play a key role as the front-end interface with customers. Under this paradigm, commercial banks would remain responsible for managing the relationship with end customers, activating the necessary know-your-customer/anti-money-laundering (KYC/AML) procedures, distributing approved CBDC wallets to customers, and resolving any issue that might arise. However, citizens would be free to receive their salaries, pensions, and other types of small incoming payments directly into their own wallets. At the same time, they would be able to spend their money by using an online or offline service provided by the central bank, which would be secure and guarantee the appropriate level of privacy.
@@
It is well known that today 137 countries and currency unions, representing 98% of the global GDP, are exploring CBDC solutions through ongoing investigations, studies, models, and pilot projects on the topic [AC_2025].
@@
In particular, the European Central Bank (ECB) plans to introduce a digital euro [ECB_2025] as a retail and wholesale CBDC. It will be freely available and free of charge for any digital payment to all citizens in the euro area. After an Investigation Phase that took place from October 2021 to October 2023, aimed at assessing feasibility, benefits, design options, possible risks, and issues, a two-year Preparation Phase was launched in November 2023. The objective is to test the potential issuance, define the necessary rules, identify the infrastructure required for smooth and secure usage, and establish the legal frameworks and operational rulebook under which to pursue the issuance. Another key element is conducting user research to ensure that the digital euro project under development is able to address the vast majority of the needs of European citizens and is inclusive from the very beginning in all its forms.
@@
From the legal point of view, in 2023 the European Commission has approved a proposal for its regulation [ECB_2025].
@@ 1.2 The Challenge of Operational Resilience in CBDC
Computing systems, along with software methodologies and programming languages needed to design and implement them, have faced a growing complexity. From sequential systems, in which a single operation at a time can be executed, the first evolution has been toward concurrent and distributed systems [Mul93], in which multiple computing devices run simultaneously and respectively exchange information through a shared memory or via message passing. The second evolution, started with the financial crisis of 2008 and the consequent decreasing trust in institutional intermediaries, has led to decentralized systems [Nak08], where parties not knowing each other perform transactions by relying on a distributed ledger and a consensus mechanism that ensure trust in an algorithmic way.
Several advanced paradigms are by now in place. They range from mobile computing, where devices are no longer tied to physical locations (IoT – Internet of Things [Mil20]), global computing, which implements the abstraction of a single computer accessible anywhere anytime (cloud, edge, fog infrastructures [EB23, BN19]), and autonomic computing, typical of systems capable of adapting to unpredictable changes (via sensors, actuators, policies [LMD13]), to machine learning, deep learning, and large language models (AI – Artificial Intelligence [RN20, WBK20, KKSS24]). Moreover, ongoing research is focusing on future paradigms such as quantum computing [NC10], which should make computation faster, and reversible computing [Lan61], which should reduce energy consumption.
@@
This high complexity exacerbate the problem of software quality; the interested reader is referred to [WikiSwBugs] for a list of major software disasters and their consequences in terms of costs, human lives, and environmental impact. Software quality is not only about code functional correctness, because it is equally important to avoid poor performance, security breaches, and bad interfaces. Moreover, the usually adopted approach of testing software is not enough as it does not guarantee the absence of errors. Therefore, it is of paramount importance to adopt software verification techniques, which are mostly based on the early development of software models and hence need to be accompanied by model-driven software engineering methodologies.
@@
The trend of cybercrimes involving hackers stealing money is closely monitored by specialized institutions. For example, the International Monetary Fund (IMF) reports more than 20,000 cyber incidents in the traditional financial sector between 2004 and 2023, with total losses of 10 billion USD (see Figure 3.2 of [IMF_2024]). In cryptocurrencies and the decentralized finance (DeFi) sector, both supported by public blockchains, Chainalysis reports nearly 14 billion USD in losses from 1,329 hacking events between 2015 and 2024 [Chainalysis_2025]. Even if not all the breaches described above are necessarily linked to structural flaws in the software, the phenomenon remains deeply concerning.
@@ 1.3 Formal Methods for Ensuring Software Quality of CBDC
To overcome these two highly undesirable situations, one might consider building two test programs. The first one, $T_{l}$, is designed so that, when supplied with the code lines of $P$, it can answer “Yes” or “No” in finite time to the question: “Does $P$ enter an infinite loop in some situations?”. The second one, $T_{f}$, is able to answer “Yes” or “No” in finite time to the question: “Does $P$ compute the function $f_{P}$?”. Unfortunately, two fundamental results of computability theory – due to Turing [Tur36] and Rice [Ric53] – establish the impossibility of constructing such tests for all programs. The practical consequence of these two results is that upon every software release there remains a residual risk of malfunction.
@@
On the specification side, formal models of software can be provided at different levels of abstraction and revolve around automata, algebras, and logics. An automaton [HMU06] consists of a set of states and a set of transitions between states. Every state of an automaton represents a state of the computation, like for instance the current statement to be executed in a program along with the value contained in every program variable, while a transition describes a state change, which can be accompanied by possible inputs and outputs. There is a hierarchy of automata classes, ranging from finite-state automata to Turing machines, which correspond to the hierarchy of formal languages established by Chomsky [Cho59]. Moreover, in addition to traditional automata in which states are global, there are variants called Petri nets [Pet62, Rei85] in which the underlying graph is bipartite so that states can be represented in a distributed way across loci of computations, which is well suited for concurrent and distributed software.
@@
Process algebras [Mil89a, Hoa85, Hen88a, BW90, BPS01] are more abstract than automata because they provide a number of operators whereby to obtain complex system models by combining simpler ones. Among these operators we mention sequential, alternative, and parallel compositions, which allow one to express the fact that two processes – intended as the behaviors of two systems – are executed one after the other or alternative to each other or run in parallel, respectively. An important notion in this setting is that of behavioral equivalence, with bisimilarity – intended as the capability of mimicking each other’s behavior stepwise – being one of the most important approaches, which identifies syntactically different processes that feature the same observable behavior. Compositional modeling is accompanied by compositional reasoning when the considered equivalence is a congruence – i.e., it allows to replace equals with equals like in arithmetic – with respect to the process algebraic operators.
@@
Logics are even more abstract in that they express the behavioral properties that a computing system should possess. The simplest example is given by Hoare logic [Hoa69, Dij75] and its variants, including separation logic [Ohe19], which are used to enrich programs with annotations containing logical formulas that state properties that should be valid in those points of programs execution. Other examples are modal and temporal logics [HC77, GHR94]. They respectively are extensions of classical logic with modal or temporal operators. The former denote the possibility or the necessity of performing certain activities in a given order. The latter express a property to be valid in the next step or the fact that a certain property has to hold until another one becomes satisfied, along at least one computation or all computations.
@@
On the verification side, we mention in particular model checking [Mcm93, CGP99, BK08] and equivalence checking [BPS01]. A model checker takes as inputs a program model, described as an automaton or a process term, and a property it should possess, formalized as a modal or temporal logic formula, then communicates whether the program model meets that property or not. In case of failure, it provides diagnostic information in the form of a program model execution that does not fulfill the property. An equivalence checker considers instead a model of the specification of the program and a model of the implementation of the program, then checks whether the two models are behaviorally equivalent to each other or not. In case of failure, it exhibits a distinguishing modal or temporal logic formula, i.e., a formula that is satisfied by only one of the two models.
@@
The aforementioned modeling and verification techniques, originally developed for functional features of programs, have been subsequently extended to deal with software architectures and quantitative aspects too [ABC10, Han92, Seg95a, ABCDF95, Hil96, Her02, AILS07, DKV09, GM14]. In this way, it is possible to address also systems including probabilistic behaviors – think, e.g., of cryptocurrencies based on proof-of-stake consensus – or real-time constraints – think, e.g., of the deadline by which a bank transfer can be canceled. Moreover, they have been adapted to address security properties as well, like for instance the use of equivalence checking to assess noninterference of information flows among different security levels [GM82, FG01, EABR25]. Many software tools have been developed – and employed in large-scale case studies – to implement formal modeling and verification, like for instance CADP [GLMS13], mCRL2 [GM14], GreatSPN [ABCDF95], PEPA Workbench [Hil96], TwoTowers [ABC10], Spin [Hol03], NuSMV [CCGR00], Uppaal [LPY97], PRISM [KNP11], Modest Toolset [HHHK13].
@@ 2 Issues and Risks of Offline Payments in a CBDC
Another important advantage of the offline payments solution is that it would enable the financial inclusion of unbanked people, who could use offline payments as a substitute for cash. This could also address the consequences of the declining use of cash [BoE_2020], with the CBDC acting as a medium for pseudonymous or anonymous payments [EUBOF_2021]. Furthermore, in a scenario in which cash is gradually abandoned as a medium for transferring value, commercial banks might find it economically disadvantageous to offer services to unbanked people, who are de facto financially excluded groups. Therefore, the introduction of offline payments as an instrument of financial inclusion is becoming a necessity to support the most vulnerable parts of our societies, which would otherwise be excluded from the economy. To achieve this, only a minimal technological infrastructure would be required – essentially, a smartphone, a basic computer, or a laptop.
In a report by the World Bank Group, it is estimated that globally 1.4 billion adults are unbanked [WBG_2021]. As for the EU, a 2024 Economic Bulletin of the ECB [ECB_2024] states that “nearly one out of five adults (19.4%) in the euro area reports not having either debit or credit cards or payment accounts”. As a consequence, the offline payments solution with cash-like features is strongly recommended as an instrument for financial inclusion also in the context of the future digital euro. Indeed, it constitutes the second requirement in the “Report on a digital euro” by the ECB [ECB_2020]:
@@
A further element supporting the offline payments solution derives from a survey conducted by the Innovation Hub of the Bank for International Settlements (BIS). It shows that “49% of central banks surveyed consider offline payments with retail CBDC to be vital, while another 49% deemed it to be advantageous[BIS_2023_Polaris_1].
Unfortunately, even if highly recommended, the offline payments solution presents some challenges and issues from operational and technical points of view. In an expository study by the Bank of Finland, a case study named “Project Pluto” is discussed, aimed at studying the risks of offline CBDCs [BoF_2023]. In contrast to the Riksbank’s e-krona pilot in Sweden [BoS_2017_Ekrona_1, BoS_2018_Ekrona_2, BoS_2022_Ekrona_3], the Pluto application focuses on the token-based model and relies on reusable tokens.
In a note of the Bank of Canada [BoC_2023] it is explicitly written that:
@@
The Bank of England also aimed to assess whether it is technically feasible to implement a secure offline payment functionality for a digital pound. The conclusions highlight some critical issues [BoE_2020]:
@@
But the most detailed study on the secure feasibility of the offline payments solution is the one conducted by BIS, which has recently produced a series of analytical documents to illustrate “Project Polaris”. The first report, entitled “A Handbook for Offline Payments with CBDC” [BIS_2023_Polaris_1], outlines the current technical solutions for offline payments with a CBDC, the possible risks, threats, and vulnerabilities associated with them, and the countermeasures to mitigate these issues. The second report, “A Security and Resilience Framework for CBDC Systems” [BIS_2023_Polaris_2], is devoted to a study of the objectives and design criteria required to achieve security and resilience in a CBDC framework. The third report, “Closing the CBDC Cyber Threat Modeling Gaps” [BIS_2023_Polaris_3], addresses the problem of analyzing the adequacy of current cybersecurity standards and frameworks through real-world observations of security flaws; the goal is to develop cyber threat models that need to be addressed to achieve full control over the security necessary in a CBDC framework. The fourth and final report, entitled “A High-Level Design Guide for Offline Payments with CBDC” [BIS_2023_Polaris_4], outlines the suggested design choices for offline payment solutions, also analyzing the security, risk, and impact on resilience of the offline system.
@@
Generally speaking, as noted in [BZKWCC_2024], we must state that basing part of the security of offline solutions on the assumption of inviolability of secure elements [^1] A secure element is a microprocessor chip with enhanced security protections that should prevent unauthorized access. It can be embedded in smart cards and smartphones or in removable subscriber identity module (SIM) cards. is simply naive. Furthermore, in general even a single compromised device could print and spend counterfeit money without restrictions. For example, a hacker able to break a single secure element could send counterfeit money to a colluding device. Moreover, without adequate countermeasures, offline payment recipients cannot distinguish legitimate payments from counterfeit, so central banks would be forced to bear these risks of fraud.
We briefly illustrate here a tentative and non-exhaustive list of threats based in part on the first report of “Project Polaris” [BIS_2023_Polaris_1] by the BIS and on the expository study of the Bank of Finland [BoF_2023]:
@@
The physical devices used in the framework of offline payments are digital electronic devices that can leak electromagnetic fields while operating. It is therefore possible, for a bad actor, to analyze the variations of these fields and perform signal analysis on the electromagnetic radiation. This could allow them to deduce some information about the electrical quantities that physically represent the bit strings associated with the digital representation of the CBDC tokens or the private key involved in a transaction.
@@
The double-spending issue can be associated with several previously described situations or contexts. Moreover, note that the double-spending issue for offline CBDC payments is different from the more famous problem of double-spending cryptocurrencies on a public and disintermediated blockchain. In the latter case, the problem of consensus on an asynchronous network must be solved, but it is well known that the extended version of the Byzantine generals problem does not admit theoretical solutions [FLP85]. Nevertheless, practical solutions based on proof-of-work [Nak08] and proof-of-stake [KN12] protocols offer some level of security. In contrast, in the case of offline payments for a CBDC, trust needs to be ensured by the tamper-resistant nature of the user device and the shared cryptographic protocols of the wallet. It is therefore necessary to ensure that double-spending attempts will be rejected by the payee’s wallet.
@@
Figure 1: The risk of devices staying offline: the bordered ones always remain offline [BoF_2023].
@@
This is probably the most insidious problem that, to the best of our knowledge, has not yet been solved. Suppose that there are two chains of offline transactions, #1 and #2, as in Figure 1. In both transaction chains, a certain amount X of tokens is moved, while the bordered smartphones always remain offline [BoF_2023].